# On the Dirichlet problem for the degenerate $k$-Hessian equation

### v1

@@ 6. The boundary estimate in terms of the interior one
$\tau=\tau(x):=\begin{pmatrix}1\\ 0\\ \vdots\\ 0\end{pmatrix}+\begin{pmatrix}~&~&~&-\rho_{11}(0)\\ ~&0_{(n-1)\times(n-1)}\hskip-28.45274pt&~&\vdots\\ ~&~&~&-\rho_{1,n-1}(0)\\ \rho_{11}(0)\hskip-19.91692pt&\cdots\hskip-17.07164pt&\rho_{1,n-1}(0)\hskip-8.53581pt&0\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\\ \vdots\\ x_{n}\end{pmatrix}=\begin{pmatrix}1-\rho_{11}(0)x_{n}\\ -\rho_{12}(0)x_{n}\\ \vdots\\ -\rho_{1,n-1}(0)x_{n}\\ \sum_{i=1}^{n-1}\rho_{1i}(0)x_{i}\end{pmatrix}$ (6.4)
@@
Without loss of generality, we assume $M\gg 1$, and $C_{\varepsilon}>1$ for all $\varepsilon>0$ in (1.8).
@@
$\langle\tau,\nu\rangle=\left[\left(-\rho_{1}\big(x^{\prime}\big)+\sum_{i=1}^{n-1}\rho_{1i}(0)x_{i}\right)+x_{n}\sum_{i=1}^{n-1}\rho_{i}\big(x^{\prime}\big)\rho_{1i}(0)\right]\frac{1}{\sqrt{1+\sum_{i=1}^{n-1}\rho_{i}^{2}}}=O\Big(\big|x^{\prime}\big|^{2}\Big)$ (6.6)
@@
$=\nabla_{x^{\prime}}\partial_{x_{i}}u\big(tx^{\prime},\rho\big(tx^{\prime}\big)\big)\cdot x^{\prime}+\partial_{x_{n}x_{i}}u\big(tx^{\prime},\rho\big(tx^{\prime}\big)\big)\left(\nabla_{x^{\prime}}\rho\big(tx^{\prime}\big)\cdot x^{\prime}\right)$
@@
$|u_{\nu}(x)-u_{n}(x)|$ $=|\nabla u(x)\cdot(\nu(x)-e_{n})|$
$\leq\sum_{i=1}^{n-1}\left|\left(u_{i}(x)-u_{i}(0)\right)\rho_{i}\big(x^{\prime}\big)\right|+|u_{n}(x)|\frac{\sum_{i=1}^{n-1}\rho_{i}^{2}(x^{\prime})}{1+\sqrt{1+\sum_{i=1}^{n-1}\rho_{i}^{2}(x^{\prime})}}$
@@
$=u_{\xi\xi}+O\Big(M\big|x^{\prime}\big|^{4}\Big)+O\Big(\big|x^{\prime}\big|^{2}\Big)+u_{i}\tau_{j}\partial_{x_{j}}(\tau_{i})$
$=u_{\xi\xi}+u_{i}\partial_{x_{1}}(\tau_{i})+u_{i}\tau_{n}\partial_{x_{n}}(\tau_{i})+O\Big(\big|x^{\prime}\big|^{2}\Big)+O\Big(M\big|x^{\prime}\big|^{4}\Big)$
$=u_{\xi\xi}+\rho_{11}(0)u_{n}-\sum_{i=1}^{n-1}\bigg(\rho_{1i}(0)\sum_{k=1}^{n-1}\rho_{1k}(0)u_{k}(x)\bigg)x_{i}+O\Big(\big|x^{\prime}\big|^{2}\Big)+O\Big(M\big|x^{\prime}\big|^{4}\Big)$
$=u_{\xi\xi}+\rho_{11}(0)u_{n}(x)+h_{1}\big(x^{\prime}\big)+O\Big(\big|x^{\prime}\big|^{2}\Big)+O\Big(M\big|x^{\prime}\big|^{4}\Big)$ (6.9)
@@
$=\left[\varphi_{11}(0)+\rho_{11}(0)\varphi_{n}(0)\right]+h_{2}\big(x^{\prime}\big)+O\Big(\big|x^{\prime}\big|^{2}\Big)-\kappa(x,\xi(x))u_{\nu}(x)$
$=\left[u_{11}(0)+\rho_{11}(0)u_{n}(0)\right]+h_{2}\big(x^{\prime}\big)+O\Big(\big|x^{\prime}\big|^{2}\Big)-\kappa(x,\xi(x))u_{\nu}(x)$ (6.10)
@@
$w(x)$ $=u_{(\tau)(\tau)}(x)-\left[u_{11}(0)+\rho_{11}(0)u_{n}(0)\right]$
$=\rho_{11}(0)u_{n}(x)-\kappa(x,\xi(x))u_{\nu}(x)+h_{1}+h_{2}+O\Big(\big|x^{\prime}\big|^{2}\Big)+O\Big(M\big|x^{\prime}\big|^{4}\Big)$
$=\left[\rho_{11}(0)-\kappa(x,\xi(x))\right]u_{n}(x)+h_{1}+h_{2}+O\Big(\big|x^{\prime}\big|^{2}\Big)+O\Big(M\big|x^{\prime}\big|^{4}\Big)$
$=\left[\rho_{11}(0)-\kappa(x,\xi(x))\right]u_{n}(0)+h_{1}+h_{2}+O\Big(\big|x^{\prime}\big|^{2}\Big)+O\Big(M\big|x^{\prime}\big|^{4}\Big)$
$=h_{3}\big(x^{\prime}\big)+h_{1}+h_{2}+O\Big(\big|x^{\prime}\big|^{2}\Big)+O\Big(M\big|x^{\prime}\big|^{4}\Big),$
@@
for all $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, where $h(x^{\prime}):=h_{1}(x^{\prime})+h_{2}(x^{\prime})+h_{3}(x^{\prime})$; $K_{0}$ and the coefficients of $h$ depend on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\|u\|_{C^{1}(\overline{\Omega})}$, second-order tangential and tangential-normal derivatives of $u$ on $\partial\Omega$. Using Theorem 1.1, we complete the proof of Lemma 6.2. ∎
@@
6.1. The best possible condition
In this subsection, we assume either $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ holds or $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$ hold. Additionally, we assume that the condition (1.5) is satisfied. Recall that $g=f^{1/p}$, where $p$ is given by (2.9).
@@ 6.1. The best possible condition
$v:=\big(x_{n}-\rho\big(x^{\prime}\big)\big)^{2}-\beta_{1}\big(x_{n}-\rho\big(x^{\prime}\big)\big)-\frac{K_{0}}{M}\big|x^{\prime}\big|^{2}-K_{0}\big|x^{\prime}\big|^{4}\quad\text{in}\ \omega_{r_{0}},$
where the constant $\beta_{1}>0$ is to be determined. It holds that
$D^{2}v=D^{2}\big[x_{n}^{2}+\beta_{1}\rho\big(x^{\prime}\big)\big]+D^{2}\big[\rho^{2}\big(x^{\prime}\big)-2x_{n}\rho\big(x^{\prime}\big)\big]-\frac{K_{0}}{M}D^{2}\Big[\big|x^{\prime}\big|^{2}\Big]-D^{2}\Big[K_{0}\big|x^{\prime}\big|^{4}\Big].$
@@
$\left\{\lambda\big(D^{2}v-\delta_{2}\mathbf{I}_{n}\big):\ x\in\omega_{r_{2}}\right\}\ \ \text{is a perturbation of}\ \ \lambda\big(D^{2}\big[x_{n}^{2}+\beta_{1}\rho\big](0)\big),\quad\forall M\geq C_{1},$
and is thus contained in a compact subset of $\Gamma_{k}$ for all $M\geq C_{1}$. We emphasize that $\beta_{1}$, $r_{2}$, $\delta_{2}$, $C_{1}$ all depend only on $\Omega$. Using the concavity of $F$, we have
$F^{ij}\big[D^{2}u\big]\left(C_{2}(v_{ij}-\delta_{2}\delta_{ij})-u_{ij}\right)$ $\geq F\big[C_{2}\big(D^{2}v-\delta_{2}\mathbf{I}_{n}\big)\big]-F\big[D^{2}u\big]$
$=C_{2}\sigma_{k}^{\frac{1}{k}}\big[D^{2}v-\delta_{2}\mathbf{I}_{n}\big]-f^{\frac{1}{k}}$
$>0$
@@
It follows from Theorems 1.1 and 1.2 that
@@
where $K_{\sigma}$ depends on $\sigma$, $\delta_{0}$, $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
@@
$K(rM+1)v=K\beta_{2}(rM+1)\big|x^{\prime}\big|^{2}\geq K\beta_{2}\left(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\right)\geq w-h\quad\text{on}\ \partial_{1}\omega_{r},$ (6.16)
by taking $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. It follows from Lemma 6.4 that
@@
$w-h\leq\left|u_{ij}\tau_{i}\tau_{j}\right|+C_{1}\leq 2\sup_{\Omega_{r^{4}}}\left|D^{2}u\right|+C_{1}$ $\leq\frac{2}{r^{4}}\left(r^{9}\sup_{\partial\Omega}\left|D^{2}u\right|+C_{r^{9}}\right)+C_{1}$
@@
$K\left(rM+\frac{C_{r^{9}}}{r^{8}}\right)v\geq K\left(rM+\frac{C_{r^{9}}}{r^{8}}\right)\frac{\beta_{1}}{2}r^{4}=\frac{K\beta_{1}}{2}Mr^{5}+\frac{K\beta_{1}}{2}\frac{C_{r^{9}}}{r^{4}}\geq w-h$ (6.18)
on $\partial_{2}\omega_{r}$, by taking the constant $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{1/(k-k/3-2/3)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
@@
$\left(u_{(\tau)(\tau)}\right)_{n}(0)$ $=\lim_{x_{n}\rightarrow 0^{+}}\frac{u_{(\tau)(\tau)}(0,x_{n})-u_{(\tau)(\tau)}(0,0)}{x_{n}}$
$=\lim_{x_{n}\rightarrow 0^{+}}\frac{w(0,x_{n})-h(0)}{x_{n}}$
$\leq K\beta_{1}\left(rM+\frac{C_{r^{9}}}{r^{8}}\right)$
@@
The proof of Lemma 6.6 is deferred to Appendix. We now proceed to prove Theorem 1.6.
@@
Without loss of generality, let the maximum $M$ be attained at the origin $0\in\partial\Omega$, $e_{n}=\nu(0)$, (6.1) hold, $\varphi(0)=0$, and $\nabla\varphi(0)=0$.
@@
where $t\in(0,1)$ depends on $x$, and the tangential vector field $\eta$ is defined in (6.3). Since $u=\varphi$ on $\partial\Omega$, the third-order tangential derivatives of $u$ are bounded (see Lemma 8.1 in the Appendix). Using Lemma 6.6, we derive
@@
for all $y\neq 0$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ and for all $\sigma>0$, where by Theorem 1.1 the constant $C_{1}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. It follows from (6.19) and (6.1) that
@@
where $C_{2}$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. Define
@@
where the constant $K>0$ is to be determined. By (6.22) and (6.23), we obtain
@@
by taking the constant $K$ large, which depends on $\delta_{0}$, $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
Using (6.21), one has
@@
by taking $K$ large, which depends on $\Omega$. It follows that
@@
on $\partial_{2}\omega_{r_{1}}\cup\partial_{3}\omega_{r_{1}}$. According to Theorem 1.1, we get
@@
where the constant $K$ depends on $\delta_{0}$, $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
@@
where $K_{\sigma}$ is given in Lemma 6.5. Taking $\sigma=1/(2K\beta_{1})$, we obtain the upper bound of $M$. This completes the proof of Theorem 1.6. ∎
6.2. The case of exponent $\mathbf{3/(2k)}$

### v2

@@ 6. The boundary estimate in terms of the interior one
$\tau=\tau(x):=\begin{pmatrix}1\\ 0\\ \vdots\\ 0\end{pmatrix}+\begin{pmatrix}~&~&~&-\rho_{11}(0)\\ ~&0_{(n-1)\times(n-1)}\hskip-28.45274pt&~&\vdots\\ ~&~&~&-\rho_{1,n-1}(0)\\ \rho_{11}(0)\hskip-19.91692pt&\cdots\hskip-17.07164pt&\rho_{1,n-1}(0)\hskip-8.53581pt&0\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\\ \vdots\\ x_{n}\end{pmatrix}$ (6.4)
@@
Without loss of generality, we assume $M\gg 1$, and $C_{\varepsilon}>1$ for all $\varepsilon$ in (1.6).
@@
$\langle\tau,\nu\rangle=O\Big(\big|x^{\prime}\big|^{2}\Big)$ (6.6)
@@
$|u_{\nu}(x)-u_{n}(x)|$ $\leq\sum_{i=1}^{n-1}\left|\left(u_{i}(x)-u_{i}(0)\right)\rho_{i}\big(x^{\prime}\big)\right|+\frac{|u_{n}(x)|\sum_{i=1}^{n-1}\rho_{i}^{2}(x^{\prime})}{1+\sqrt{1+\sum_{i=1}^{n-1}\rho_{i}^{2}}}$
@@
$=u_{\xi\xi}+u_{i}\tau_{j}\partial_{x_{j}}(\tau_{i})+O\Big(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\Big)$
$=u_{\xi\xi}+\rho_{11}(0)u_{n}(x)+h_{1}\big(x^{\prime}\big)+O\Big(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\Big)$ (6.9)
@@
$=[\varphi_{11}+\rho_{11}\varphi_{n}](0)+h_{2}\big(x^{\prime}\big)+O\Big(\big|x^{\prime}\big|^{2}\Big)-\kappa(x,\xi(x))u_{\nu}(x)$
$=[u_{11}+\rho_{11}u_{n}](0)+h_{2}\big(x^{\prime}\big)+O\Big(\big|x^{\prime}\big|^{2}\Big)-\kappa(x,\xi(x))u_{\nu}(x)$ (6.10)
@@
$w(x)$ $=u_{(\tau)(\tau)}(x)-[u_{11}+\rho_{11}u_{n}](0)$
$=\rho_{11}(0)u_{n}(x)-\kappa(x,\xi(x))u_{\nu}(x)+h_{1}+h_{2}+O\Big(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\Big)$
$=\left[\rho_{11}(0)-\kappa(x,\xi(x))\right]u_{n}(x)+h_{1}+h_{2}+O\Big(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\Big)$
$=\left[\rho_{11}(0)-\kappa(x,\xi(x))\right]u_{n}(0)+h_{1}+h_{2}+O\Big(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\Big)$
$=h_{3}\big(x^{\prime}\big)+h_{1}+h_{2}+O\Big(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\Big),$
@@
for all $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, where $h(x^{\prime}):=h_{1}(x^{\prime})+h_{2}(x^{\prime})+h_{3}(x^{\prime})$; $K_{0}$ and the coefficients of $h$ depend on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\|u\|_{C^{1}(\overline{\Omega})}$, second-order tangential and tangential-normal derivatives of $u$ on $\partial\Omega$. Using the results in Section 3, we complete the proof of Lemma 6.2. ∎
@@
6.1. The case under condition $(i)$ or $(ii)$
In this subsection, we assume condition (1.7) is satisfied, and either $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ holds or $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$ hold.
@@ 6.1. The case under condition $(i)$ or $(ii)$
$v:=\big(x_{n}-\rho\big(x^{\prime}\big)\big)^{2}-\beta_{1}\big(x_{n}-\rho\big(x^{\prime}\big)\big)-\frac{K_{0}}{M}\big|x^{\prime}\big|^{2}-K_{0}\big|x^{\prime}\big|^{4}$
in $\omega_{r_{0}}$, where the constant $\beta_{1}>0$ is to be determined. It holds that
$D^{2}v$ $=D^{2}\big[x_{n}^{2}+\beta_{1}\rho\big(x^{\prime}\big)\big]+D^{2}\big[\rho^{2}\big(x^{\prime}\big)-2x_{n}\rho\big(x^{\prime}\big)\big]$
$\hskip 12.80373pt-\frac{K_{0}}{M}D^{2}\Big[\big|x^{\prime}\big|^{2}\Big]-D^{2}\Big[K_{0}\big|x^{\prime}\big|^{4}\Big].$
@@
$\left\{\lambda\big(D^{2}v-\delta_{2}\mathbf{I}_{n}\big):\ x\in\omega_{r_{2}}\right\}\ \ \text{is a perturbation of}\ \ \lambda\big(D^{2}\big[x_{n}^{2}+\beta_{1}\rho\big](0)\big)$
for all $M\geq C_{1}$, and is thus contained in a compact subset of $\Gamma_{k}$ for all $M\geq C_{1}$. We emphasize that $\beta_{1}$, $r_{2}$, $\delta_{2}$, $C_{1}$ all depend only on $\Omega$. Using the concavity of $F$, we have
$F^{ij}\big[D^{2}u\big]\left(C_{2}(v_{ij}-\delta_{2}\delta_{ij})-u_{ij}\right)\geq C_{2}\sigma_{k}^{\frac{1}{k}}\big[D^{2}v-\delta_{2}\mathbf{I}_{n}\big]-f^{\frac{1}{k}}>0$
@@
It follows from the results in Section 3 that
@@
where constant $K_{\sigma}$ depends on $\sigma$, $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
@@
where constant $K$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$.
@@
$K(rM+1)v=K\beta_{2}(rM+1)\big|x^{\prime}\big|^{2}\geq K\beta_{2}\left(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\right)\geq w-h$ (6.16)
on $\partial_{1}\omega_{r}$, by taking $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. It follows from Lemma 6.4 that
@@
$w-h\leq\left|u_{ij}\tau_{i}\tau_{j}\right|+C_{1}$ $\leq 2\sup_{\Omega_{r^{4}}}\left|D^{2}u\right|+C_{1}$
$\leq\frac{2}{r^{4}}\left(r^{9}\sup_{\partial\Omega}\left|D^{2}u\right|+C_{r^{9}}\right)+C_{1}$
@@
$-K\left(rM+\frac{C_{r^{9}}}{r^{8}}\right)v\geq K\left(rM+\frac{C_{r^{9}}}{r^{8}}\right)\frac{\beta_{1}}{2}r^{4}$ $=\frac{K\beta_{1}}{2}Mr^{5}+\frac{K\beta_{1}}{2}\frac{C_{r^{9}}}{r^{4}}$
$\geq w-h$ (6.18)
on $\partial_{2}\omega_{r}$, by taking $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
@@
$\left(u_{(\tau)(\tau)}\right)_{n}(0)=\lim_{x_{n}\rightarrow 0^{+}}\frac{w(0,x_{n})-h(0)}{x_{n}}\leq K\beta_{1}\left(rM+\frac{C_{r^{9}}}{r^{8}}\right)$
@@
Proof.
It is clear that the expression $(u_{\nu ij}\eta_{i}\eta_{j}+u_{\nu i}\eta_{j}\partial_{x_{j}}(\eta_{i}))(x)$ is invariant under orthogonal transformations.
Step 1. Let $\tilde{\eta}(x)$ be an arbitrary tangential vector field defined in $B_{r}(0)$, where the radius $r>0$ can be arbitrarily small depending on the origin $0\in\partial\Omega$. Furthermore, let $\tilde{\eta}$ satisfy
for some constant $C$.
Differentiating $\nu\cdot\tilde{\eta}(x^{\prime},\rho(x^{\prime}))=0$ with respect to $x_{1}$ at the origin yields
$\partial_{x_{1}}(\tilde{\eta}_{n})(0)-\rho_{11}(0)=0.$
Then we have
$u_{n(\tilde{\eta})(\tilde{\eta})}(0)=u_{n11}(0)+u_{ni}(0)\partial_{x_{1}}(\tilde{\eta}_{i})(0)\leq u_{n11}(0)+u_{nn}(0)\rho_{11}(0)+K_{1},$
where $K_{1}$ depends on $C$ and tangential-normal derivatives of $u$ on $\partial\Omega$. It holds that
$\left(u_{(\tau)(\tau)}\right)_{n}(0)$ $=u_{n11}(0)+u_{in}(0)\tau_{j}\partial_{x_{j}}(\tau_{i})(0)$
$\hskip 12.80373pt+u_{ij}\partial_{x_{n}}[\tau_{i}\tau_{j}]+u_{i}\partial_{x_{n}}\big[\tau_{j}\partial_{x_{j}}(\tau_{i})\big]$
$\geq u_{n11}(0)+u_{nn}(0)\rho_{11}(0)-K_{2},$
$u_{n(\tilde{\eta})(\tilde{\eta})}(0)\leq\left(u_{(\tau)(\tau)}\right)_{n}(0)+K_{1}+K_{2}\leq\sigma M+K_{\sigma},\quad\forall\sigma>0,$ (6.19)
by Lemma 6.5 and increasing $K_{\sigma}$.
Step 2. For any $x_{0}\neq 0$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, by rotation and translation, we introduce a new coordinate system $y$ such that $x_{0}$ is the origin $0\in\partial\Omega$, $e_{n}=\nu(0)$, and $\eta(x_{0})$ is in the direction $e_{1}$. Denote the expression of vector field $\eta(x)$ in the new coordinate system by $\bar{\eta}(y)$. It is clear that $\bar{\eta}$ is well-defined in $B_{|x_{0}|/2}(0)$. Recalling (6.3), by calculation, one can verify
$\left|\partial_{y_{1}}(\bar{\eta}_{i})(0)\right|\leq C,\quad\forall 1\leq i\leq n-1,$
where $C$ depends only on the $C^{2}$-norm of $\partial\Omega$.
Define $\tilde{\eta}(y):=\bar{\eta}(y)/|\eta(x_{0})|$. Then
$\tilde{\eta}(0)=e_{1}\quad\text{and}\quad\left|\partial_{y_{1}}(\tilde{\eta}_{i})(0)\right|\leq C,\quad\forall 1\leq i\leq n-1.$
By (6.19) and $|\eta(x_{0})|\leq 2$, we get
$\left(u_{nij}\bar{\eta}_{i}\bar{\eta}_{j}+u_{ni}\bar{\eta}_{j}\partial_{y_{j}}(\bar{\eta}_{i})\right)(0)\leq 4u_{n(\tilde{\eta})(\tilde{\eta})}(0)\leq\sigma M+4K_{\sigma/4}$
for all $\sigma>0$. Coming back to the original coordinate system (6.1), by increasing $K_{\sigma}$, we derive
for all $x_{0}\neq 0$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$. The proof of Lemma 6.6 is complete. ∎
We now proceed to establish the boundary second-order normal derivative estimate.
@@
Without loss of generality, let $M$ be attained at the origin $0\in\partial\Omega$, $e_{n}=\nu(0)$, (6.1) hold, $\varphi(0)=0$, and $\nabla\varphi(0)=0$.
@@
where $t\in(0,1)$ depends on $x$, and $\eta$ is defined in (6.3). Differentiating $u=\varphi$ on $\partial\Omega$ three times, yields that the third-order tangential derivatives of $u$ are bounded. Using Lemma 6.6, we derive
@@
for all $y\neq 0$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ and for all $\sigma>0$, where $C_{1}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. It follows from (6.20) and (6.1) that
@@
where the constant $C_{2}$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. Define
@@
where $K>0$ is a constant to be determined. By (6.23) and (6.24), we obtain
@@
by taking the constant $K$ large, which depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
The boundary $\partial\omega_{r_{1}}$ consists of three parts: $\partial_{1}\omega_{r_{1}}\cup\partial_{2}\omega_{r_{1}}\cup\partial_{3}\omega_{r_{1}}$, defined in the proof of Lemma 3.2. Using (6.22), one has
@@
by taking $K$ large, which depends on $\Omega$. It is clear that
@@
on $\partial_{2}\omega_{r_{1}}\cup\partial_{3}\omega_{r_{1}}$. According to the results in Section 3, we get
@@
where the constant $K$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
@@
where $K_{\sigma}$ is given in Lemma 6.5. Taking $\sigma=1/(2K\beta_{1})$, we obtain the upper bound of $M$. Then the $C^{2}$ estimate independent of $\inf_{\Omega_{0}}f$ is established, and Theorem 1.3 follows from an approximation argument. ∎
6.2. The case under condition $(iii)$